OpenStudy (anonymous):

The volume of a square pyramid is equal to _____ times the volume of a cube where the bases are the same, but the square pyramid is half the height of the cube. A. 1/6 B. 1/3 C. 3 D. 6

OpenStudy (jack1):

Volume of a pyramid = ( L x W x H )/3 Volume of a cube = L x W X H then: 3 x Vp = ( Vc) so if L and W are the same, but Hp = 1/2 x Hc Vp = 1/6 x ( Vc)